INFRARED SPECTROSCOPY OF LARGER BENZENE-(H_{2}O)_{n} HYDROGEN-BONDED CLUSTERS: EVI-DENCE OF CUBIC STRUCTURES

\underline{C. J. GRUENLOH}, C. A. ARRINGTON, J. R. CARNEY, F. C. HAGEMEISTER, and T. S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393; S. Y. FREDERICKS and K. D. JORDAN, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

One color R2PI-TOFMS, ultraviolet hole-burning, and RIDIRS have been used to assign and characterize the hydrogenbonding topology of size-selected benzene-(H_{2}O)_{n} clusters formed in a supersonic molecular beam. Hole-burning spectroscopy confirms the presence of conformers in the increasingly congested R2PI spectra of these larger clusters. Two transitions in the benzene-(H_{2}O)_{6}^{+} R2PI-TOFMS are assigned to the benzene-(H_{2}O)_{8} cluster. Comparison of the experimental RIDIR spectra to density functional theory calculations results in the assignment of these conformers to cubic octamers of S_{4} and D_{2d} symmetry. The presence of benzene weakly distorts the cubic structures, breaking the degeneracy of several OH stretch vibrations and inducing intensity in otherwise forbidden infrared transitions. Four transitions in the benzene-(H_{2}O)_{7}^{+} R2PI-TOFMS are tentatively assigned as two conformers each of the (benzene)_{2}-(H_{2}O)_{8} and the benzene-(H_{2}O)_{9} clusters. The RIDIRS of these species share common features with those of the cubic benzene-(H_{2}O)_{8} octamers; thus, a structure for the benzene-water nonamer built by adding one water molecule to a cubic octamer is tested against calculations.